Show all work and no excel
FORMATION r r?E ?N BELOW FOR THE FOLLOWING PROBLEMS 29-30 DPS Total Asset Turnover Net Profit Margin EPS Total Assets/Equity Wal-Blue 1.00 3.20 3.50% 4.00 3.00 Industry 1.50 2.50 3.00% 3.00 4.00 29. What is the ROE for Wal-Blue? 30. What is the expected sustainable growth rate for the industry? 31. A firm has a current price of $40 a share, an expected growth rate of 11 percent and expected dividend per share (DI) of $2. Given its risk you have a required rate of return for it of 12 percent. What is your expected rate of return and should you buy-yes/no?
Answer #1

29) ROE = Profit Margin x Asset Turnover x Asset / Equity

= 3.5% x 3.2 x 3 = 33.6%

30) Sustainable Growth Rate = ROE x (1 - payout ratio)

Industry ROE = 2.5% x 3 x 4 = 30% and payout ratio = 1 - DPS / EPS = 1 - 1.5/3 = 50%

=> Growth rate = 30% x (1 - 50%) = 15%

31) Expected Rate of return = D1 / P + g = 2 / 40 + 11% = 16%

As expected returns are higher than required returns, you should buy buy it.